OpenStudy (lizilizi28):

· Write a word problem involving a quadratic function. How would you explain the steps in finding the solution

OpenStudy (anonymous):

do you know what a quadratic is?

OpenStudy (anonymous):

5x^2-32x+12 thats a quadratic function solve it: 5x^2-30x-2x+12=0 5x(x-6) -2(x-6) (5x-2) (x-6) x=2/5 and x=6

OpenStudy (anonymous):

now making a word problem is your job :)

OpenStudy (lizilizi28):

ok than you